OpenStudy (anonymous):

secx+tanx=cosx/1−sinx Prove this.

OpenStudy (akashdeepdeb):

secx + tanx = 1/cosx + sinx/cosx = (1+sinx)/cosx = Multiply denominator and the numerator by (1-sinx) It becomes = (1-sin^2x)/cosx(1-sinx) = (cos^2x/cosx(1-sinx) = cosx/1 - sinx = RHS Understand? :)

OpenStudy (anonymous):

Ah yes, that was it! I'll have to remeber that